what will be the mass of a certain gas if it subjected to a 6 meter diameter pressurized cylindrical container at 40 degree celcius and 110 kpa the gas is initially measured from its standard temperature and pressure which calculated its density to be 4 kg/m^3 the height of the container is 6 meters


1
Expert's answer
2021-12-21T16:30:03-0500
"m=\rho V=0.25\pi d^2h\rho\\
m=0.25\pi (6)^2(6)(4)=679\ kg"